the white hot pepper is a traditional jazz band.the length in minutes of each piece of music played by the band may be modelled by a normal distribution with mean 5 and standard deviation 1.5 and may be assumed to be independent of the length of all other pieces.

find the probability that a particular piece will last between 3.5 and 7.25 minutes

find the probability that the average length of the next 6pieces to be played by the band will be less than 4 minutes.

there is an interval of exactly one minute btw the band finishing one piece and starting the next. the band starts to play its last 6pieces at 10:31pm. using your answer above state whether you think the band will still be playing at 11:00pm. justify your answer



1
Expert's answer
2021-07-03T09:15:19-0400

a. Let "X=" the length in minutes of each piece of music played by the band: "X\sim N(\mu , \sigma^2)."

Given "\mu=5, \sigma=1.5"


"P(3.5<X<7.25)=P(X<7.25)-P(X\leq3.5)"

"=P(Z<\dfrac{7.25-5}{1.5})-P(Z\leq\dfrac{3.5-5}{1.5})"

"=P(Z<1.5)-P(Z\leq-1)"

"\approx0.9331928-0.1586553"

"\approx0.7745"

b. Let "\bar{X}=" the average length i: "\bar{X}\sim N(\mu , \sigma^2/n)."

Given "\mu=5, \sigma=1.5, n=6"

"P(\bar{X}<4)=P(Z<\dfrac{4-5}{1.5/\sqrt{6}})"

"\approx P(Z<-1.633)\approx0.0512"

c. Let "Y=" the length of 6 pieces with intervals: "Y=6X+5"

"Y\sim N(\mu_1, \sigma_1^2)"


"\mu_1=6(5)+5=35"

"\sigma_1^2=(6)^2(\sigma)^2=81"

"P(Y>29)=1-P(Y\leq29)"

"=1-P(Z\leq \dfrac{29-35}{9})\approx1-P(Z\leq -0.6667)"

"\approx0.7475"

Since "0.7475>0.7," I think the band will still be playing at 11:00pm.